# Starlight from JWST: Implications for star formation and dark matter models

### v2

@@
Juan Urrutia , juan.urrutia@kbfi.ee1122    John Ellis , john.ellis@cern.ch3344    Malcolm Fairbairn , malcolm.fairbairn@kcl.ac.uk33    Ville Vaskonen , ville.vaskonen@pd.infn.it115566
We confront the star formation rate in different dark matter (DM) models with UV luminosity data from JWST up to $z\simeq 25$ and legacy data from HST. We find that a transition from a Salpeter population to top-heavy Pop-III stars is likely at $z\simeq 10$ and that beyond $z=10-15$ the feedback from supernovae and active galactic nuclei is progressively reduced, so that at $z\simeq 25$ the production of stars is almost free from any feedback. We compare fuzzy and warm DM models that suppress small-scale structures with the CDM paradigm, finding that the fuzzy DM mass $>4.5\times 10^{-22}{\rm eV}$ and the warm DM mass $>1.5\,{\rm keV}$ at the 95% CL. The fits of the star formation rate parametrization do not depend strongly on the DM properties within the allowed range. We find no preference over CDM for enhanced matter perturbations associated with axion miniclusters or primordial black holes. The scale of the enhancement of the power spectrum should be $>27\,{\rm Mpc}^{-1}$ at the 95% CL, excluding axion miniclusters produced for $m_{a}<7.5\times 10^{-17}\,{\rm eV}$ or heavy primordial black holes that constitute a fraction $f_{\rm PBH}>\max[88M_{\odot}/m_{\rm PBH},10^{-4}(m_{\rm PBH}/10^{4}M_{\odot})^{-0.09}]$ of DM.
@@ 1 Introduction
An important target for James Webb Space Telescope (JWST) observations is to measure the UV luminosity function at high $z$ (Naidu et al. 2022; Castellano et al. 2023; Donnan et al. 2023; Morishita & Stiavelli 2023; Bouwens et al. 2023; Pérez-González et al. 2023; Harikane et al. 2023, 2024; Pérez-González et al. 2025; Castellano et al. 2025). These observations complement the more precise lower-$z$ observations from the Hubble Space Telescope (HST) (McLeod et al. 2016; Oesch et al. 2018; Bouwens et al. 2021; Finkelstein et al. 2022; Leethochawalit et al. 2023; Bagley et al. 2024). The high-$z$ observations serve as a probe of galaxy dynamics and the star formation rate (SFR) at early epochs of the universe, constraining the abundance and masses of the conjectured first-generation Pop-III stars (Maiolino et al. 2024; Vanzella et al. 2023; Ventura et al. 2024; Fujimoto et al. 2025). The frontier of UV luminosity function measurements has been pushed back to unprecedented early times with recent measurements at $z\sim 25$, opening a window on the epoch before reionization took place (Pérez-González et al. 2025; Castellano et al. 2025).
@@
The number density of Pop-III stars depends on the abundance of early galaxies. This depends in turn on the nature of dark matter (DM), since galaxies populate DM halos. The UV luminosity function is, therefore, a valuable source of insights into DM models (Sabti et al. 2022b; Hütsi et al. 2023; Parashari & Laha 2023; Liu et al. 2024; Winch et al. 2024). As the redshift increases, the halo mass function (HMF) becomes more sensitive to the physical properties of DM, since fewer halos are built via hierarchical growth and the imprints from the physics before CMB are more noticeable.
@@
Our understanding of structure formation and the star formation process in models of DM with suppressed small-scale structure formation at high $z\gtrsim 10$, such as Fuzzy Dark Matter (FDM) and Warm Dark Matter (WDM), or in models with an enhanced matter power spectrum, is very limited. Therefore, the very high redshift abundance and distribution of Pop-III stars, made of pristine gas from the Big Bang, are topics of hot debate. The roles of external radiation, internal fragmentation of the gas clouds and the dynamics that form the stars are highly uncertain, so the environments where these first stars were born and their mass function are as yet unknown; see Klessen & Glover (2023) for a review. The UV luminosity function measured with JWST has proven to be higher than anticipated, even when considering only the galaxies which are spectroscopically confirmed (Harikane et al. 2024). This suggests that either the feedback mechanisms preventing star formation were less restrictive (Susa & Umemura 2004; Dekel et al. 2023; Fukushima & Yajima 2021), or that the Pop-III stars were very top-heavy and therefore more luminous (Hirano et al. 2015; Steinhardt et al. 2023; Hirano et al. 2014). This tension has been exacerbated by the measurements of starlight at $z\sim 25$, which are higher than the predictions from many semi-analytical models in the literature (Pérez-González et al. 2025).
@@
The objectives of this work are to parameterize the essential physics of high-$z$ star formation and understand the features necessary to reproduce the UV luminosity function back to $z\sim 25$ in different DM models. We take into account possible degeneracies and features of modified DM models, feedbacks and changes in the stellar population. To this end, we connect the HMF obtained using the excursion set formalism (Bond et al. 1991), with the early Universe microphysics of DM, and combine this with a phenomenological parameterization of star formation. We also discuss the effects of gravitational lensing and dust attenuation, neither of which seem to be very significant for the interpretation of the $z>15$ UV data.
@@
To extrapolate the results of DM simulations from low- to high-$z$ and get reliable bounds, we improve on existing techniques in the literature in two key aspects: 1) We derive a new estimate of the DM halo growth rate that accounts for the ellipsoidal collapse correction, and use it consistently for different DM models. To this end, we recompute the first-crossing probability distributions of the paths with a moving barrier and find a better fit than that originally proposed by Sheth et al. (2001). The conditional mass function in the Sheth-Tormen approach is divergent for very small steps, as has been pointed out by Jiang & van den Bosch (2014) and Zhang et al. (2008). Our fit is divergence-free and reproduces analytical estimations of Zhang et al. (2008). 2) We provide new values for the parameters the smooth-$k$ window function that, unlike those proposed in Leo et al. (2018), match consistently the CDM prediction.
@@
We find that, under the assumption that the star formation rate (SFR) is proportional to the halo growth rate, the observations of the UV luminosity function up to high $z$ indicate two interesting effects that change the luminosity function. First, at $z>10$, either the stars need to be more luminous and/or the SFR needs to be higher. This transition is relatively sharp and might be caused by a top-heavy Pop III-dominated star population. We find that the evidence for this effect above the $95\%$ CL. In addition, the SFR needs to be further enhanced at $z>15$. This can be achieved by reducing the feedback from supernovae (SNe) and active galactic nuclei (AGNs) on the SFR. We find that to explain the $z>15$ observations, these feedbacks need to be significantly reduced to match the UV luminosity function, and at $z\simeq 25$ the production of stars should be almost free from feedback effects. In agreement with Yung et al. (2025), we find that, with these changes in the stellar population and feedback effects, the JWST data can be explained within the CDM model.
@@
An important focus of this work is whether different DM scenarios require different star formation physics to fit high redshift data. Interestingly, we find very little sensitivity to DM properties that differ from CDM, with only a small change in star formation rates favoured for WDM, and no change for FDM.
@@
When marginalizing over the SFR parametrization, we find a lower bound of $m_{\rm FDM}>4.5\times 10^{-22}{\rm eV}$ for the case of FDM and $m_{\rm WDM}>1.5\,{\rm keV}$ for the case of WDM. The new JWST data, although pushing the observations to much earlier galaxies, are still not precise enough to put much stronger bounds than the HST data, as also found in Winch et al. (2024). We also find that enhanced small-scale matter perturbations are not preferred over changes in the star formation due to the presence of Pop-III stars and changes in the feedback in the pre-reionisation era. We are able to put bounds on the scale of white-noise enhanced perturbations at $k_{c}>27\,{\rm Mpc}^{-1}$, excluding, for example, the possibility of axion miniclusters produced for $m_{a}<7.5\times 10^{-17}\,{\rm eV}$ or heavy primordial black holes (PBHs) that constitute the fraction $f_{\rm PBH}>\max[88M_{\odot}/m_{\rm PBH},10^{-4}(m_{\rm PBH}/10^{4}M_{\odot})^{-0.09}]$ of DM. In a forthcoming paper, we will leverage these findings to compute the growth of SMBHs in non-CDM scenarios and obtain, for the first time, even more competitive bounds on deviations from CDM using the measurements of high-$z$ SMBHs by JWST.
@@
Throughout this work, we use the Planck 2018 CMB best-fit values for the cosmological parameters (Aghanim et al. 2020): $\Omega_{\rm M}=0.315$, $\Omega_{\rm B}=0.0493$, $z_{\rm eq}=3402.0$, $\sigma_{8}=0.811$, $h=0.674$, $T_{0}=2.7255\,{\rm K}$ and $n_{s}=0.965$.
@@ 2 Halo mass function and growth rate
Figure 1: Left panel: The unconditional first-crossing distribution for walks starting from $S=0$, $\delta=0$. The histogram shows the results of random walks, the solid black and dashed orange curves show the ansatz (3) with $q=0.80$ and $q=0.707$, respectively, and the dotted gray curve shows the spherical collapse result. Other panels: The conditional first-crossing distribution at $z>z_{0}$ for walks starting from $S_{0}>0$ and $\delta_{\rm ell}(z_{0},S_{0})>0$ as indicated in the plot. The histogram again shows the results of random walks, the solid black curves show the ansatz (LABEL:eq:ST2), the dashed orange curves show the ansatz (3) with $\delta_{\rm sp}(z)\to\delta_{\rm sp}(z)-\delta_{\rm sp}(z_{0})$ and $S\to S-S_{0}$, and the dotted gray curve shows the spherical collapse result.
In the excursion-set formalism (Bond et al. 1991), the conditional and unconditional HMFs are obtained from the first-threshold-crossing distributions of uncorrelated random walks in the $(S,\delta)$ plane, where $S=\sigma(M)^{2}$ denotes the variance of matter perturbations at mass scale $M$ and $\delta$ is the matter density contrast around a given spatial point. The unconditional mass function is obtained from the first-crossing distribution $p_{\rm FC}(S,z)$ of walks starting from $(0,0)$:
@@
For spherical collapse in a universe with CDM, the density contrast threshold $\delta_{\rm sp}(z)=3(3\pi/2)^{2/3}/(5D_{\rm g}(z))$ (Dodelson 2003), where the CDM linear growth function $D_{\rm g}(z)$ does not depend on $S$ and the first-crossing distribution $p_{\rm FC}(S,z)$ can be computed analytically. In the extension to include ellipsoidal collapse, the threshold becomes $S$-dependent:
@@
with $a=0.707$ (Sheth et al. 2001), and $p_{\rm FC}(S,z)$ cannot be calculated analytically. The standard Sheth-Tormen ansatz for the first-crossing distribution with ellipsoidal collapse is (Sheth & Tormen 1999)
@@
[…]
For the conditional mass function, we study random walks starting from $S_{0}>0$ and $\delta_{0}=\delta_{\rm ell}(S_{0},z_{0})>0$. The simplest ansatz for the first-crossing distribution $p_{\rm FC}(S,z|S_{0},z_{0})$, as considered, e.g., by Sheth & Tormen (2002), is obtained by the shift $\delta_{\rm sp}(z)\to\delta_{\rm sp}(z)-\delta_{\rm sp}(z_{0})$ and $S\to S-S_{0}$ in (3). As seen in Fig. 1, comparing the orange dashed curves with the histograms we find that this ansatz does not reproduce well the results of the random walks. A better estimate, shown by the solid black curves in Fig. 1, is given by
@@
[…]
@@
Unlike the estimate derived by Correa et al. (2015), our estimate (7) accounts for the correction from ellipsoidal collapse. Moreover, our estimate is based on the expected increase in the halo mass while the estimate of Correa et al. (2015) is based on finding the expected mass of the main progenitor. However, we find that these estimates agree within a factor two across a broad range of halo masses and redshifts.
@@ 3.1 Suppressed small-scale structures
where $\mathcal{P}_{\rm CDM}(k)$ denotes the CDM matter power spectrum that we compute with the transfer function from Eisenstein & Hu (1998).
In the case of FDM, the speed of sound is modified by quantum pressure, so that $c_{s}\sim k/(4a^{2}m_{\rm FDM}^{2})$. Consequently, the Jeans scale is non-zero (see e.g. Marsh (2016)),
@@
and perturbations at small scales $k\lesssim k_{\rm J}$ are suppressed. This suppression can be estimated using an effective fluid approximation that is accurately matched to the solution of the Klein-Gordon equation (Hu et al. 2000; Passaglia & Hu 2022). We use the fitting formula provided by Passaglia & Hu (2022):
@@
In the case of WDM, a thermal production mechanism is typically assumed and the particles are lighter than typical CDM particles, remaining relativistic for a longer time. This implies that the free streaming length of WDM particles can be long, $\lambda_{\rm fs}\sim(m_{\rm WDM}/{\rm keV})^{-4/3}{\rm Mpc}$. In order to compute the resultant damping of the small-scale structures, it is necessary to track accurately the transition between relativistic and non-relativistic behaviour by solving the Boltzmann equations. This yields the fit (Bode et al. 2001; Hansen et al. 2002; Viel et al. 2005)

### v4

@@
Juan Urrutia , juan.urrutia@kbfi.ee    John Ellis ,    Malcolm Fairbairn ,    Ville Vaskonen ,
We confront the star formation rate in different dark matter (DM) models with UV luminosity data from JWST up to $z\simeq 25$ and legacy data from HST. We find that a transition from a Salpeter population to top-heavy Pop-III stars is likely at $z\simeq 10$ and that beyond z=1015z=10-15 the feedback from supernovae and active galactic nuclei is progressively reduced, so that at $z\simeq 25$ the production of stars is almost free from any feedback. We compare fuzzy and warm DM models that suppress small-scale structures with the CDM paradigm, finding that the fuzzy DM mass $5.6\times 10^{-22}{\rm eV}$ and the warm DM mass $1.5\,{\rm keV}$ at the 95% CL. The fits of the star formation rate parametrisation do not depend strongly on the DM properties within the allowed range. We find no preference over CDM for enhanced matter perturbations associated with axion miniclusters or primordial black holes. The scale of the enhancement of the power spectrum should be $25\,{\rm Mpc}^{-1}$ at the 95% CL, excluding axion miniclusters produced for $m_{a}<6.6\times 10^{-17}\,{\rm eV}$ or heavy primordial black holes that constitute a fraction $f_{\rm PBH}>\max[105M_{\odot}/m_{\rm PBH},10^{-4}(m_{\rm PBH}/10^{4}M_{\odot})^{-0.09}]$ of DM.
@@ 1 Introduction
An important target for James Webb Space Telescope (JWST) observations is to measure the UV luminosity function at high $z$ (2022ApJ...940L..14N; Castellano:2022ikm; 2023MNRAS.518.6011D; 2023ApJ...946L..35M; 2023MNRAS.523.1036B; 2023ApJ...951L...1P; 2023ApJS..265....5H; 2024ApJ...960...56H; Perez-Gonzalez:2025bqr; Castellano:2025vkt). These observations complement the more precise lower-$z$ observations from the Hubble Space Telescope (HST) (2016MNRAS.459.3812M; 2018ApJ...855..105O; 2021AJ....162...47B; 2022ApJ...928...52F; 2023MNRAS.524.5454L; 2024ApJ...961..209B). The high-$z$ observations serve as a probe of galaxy dynamics and the star formation rate (SFR) at early epochs of the universe, constraining the abundance and masses of the conjectured first-generation Pop-III stars (Maiolino:2023wwm; 2023A&A...678A.173V; Ventura:2024mbp; Fujimoto:2025kbv). The frontier of UV luminosity function measurements has been pushed back to unprecedented early times with recent measurements at $z\sim 25$, opening a window on the epoch before reionisation took place (Perez-Gonzalez:2025bqr; Castellano:2025vkt).
@@
The number density of Pop-III stars depends on the abundance of early galaxies. This depends in turn on the nature of dark matter (DM), since galaxies populate DM halos. The UV luminosity function is, therefore, a valuable source of insights into DM models (Sabti:2021unj; Hutsi:2022fzw; Parashari:2023cui; Liu:2024edl; Winch:2024mrt; Sipple:2024svt). As the redshift increases, the halo mass function (HMF) becomes more sensitive to the physical properties of DM, since fewer halos are built via hierarchical growth and the imprints from the physics before the CMB are more noticeable.
@@
Our understanding of structure formation and the star formation process in models of DM with suppressed small-scale structure formation at high $z\gtrsim 10$, such as Fuzzy Dark Matter (FDM) and Warm Dark Matter (WDM), or in models with an enhanced matter power spectrum, is very limited. Therefore, the very high redshift abundance and distribution of Pop-III stars, made of pristine gas from the Big Bang, are topics of hot debate. The roles of external radiation, internal fragmentation of the gas clouds and the dynamics that form the stars are highly uncertain, so the environments where these first stars were born and their mass function are as yet unknown; see Klessen:2023qmc for a review. The UV luminosity function measured with JWST has proven to be higher than anticipated, even when considering only the galaxies which are spectroscopically confirmed (2024ApJ...960...56H). This suggests that either the feedback mechanisms preventing star formation were less restrictive (2004ApJ...600....1S; 2023MNRAS.523.3201D; 2021MNRAS.506.5512F), or that the Pop-III stars were very top-heavy and therefore more luminous (2015MNRAS.448..568H; 2023ApJ...951L..40S; 2014ApJ...781...60H). This tension has been exacerbated by the measurements of starlight at $z\sim 25$, which are higher than the predictions from many semi-analytical models in the literature (Perez-Gonzalez:2025bqr).
@@
The objectives of this work are to parameterise the essential physics of high-$z$ star formation and understand the features necessary to reproduce the UV luminosity function back to $z\sim 25$ in different DM models. We take into account possible degeneracies and features of modified DM models, feedback and changes in the stellar population. To this end, we connect the HMF obtained using the excursion set formalism (Bond:1990iw), with the early Universe microphysics of DM, and combine this with a phenomenological parameterisation of star formation. We also discuss the effects of gravitational lensing and dust attenuation, neither of which seems to be very significant for the interpretation of the $z>15$ UV data.
@@
To extrapolate the results of DM simulations from low- to high-$z$ and get reliable bounds, we improve on existing techniques in the literature in two key aspects: 1) We derive a new estimate of the DM halo growth rate that accounts for the ellipsoidal collapse correction, and use it consistently for different DM models. To this end, we recompute the first-crossing probability distributions of the paths with a moving barrier and find a better fit than that originally proposed by Sheth:1999su. The conditional mass function in the Sheth-Tormen approach is divergent for very small steps, as has been pointed out by Jiang:2013kza and Zhang:2008dx. Our fit is divergence-free and reproduces analytical estimations of Zhang:2008dx. 2) We provide new values for the parameters the smooth-$k$ window function that, unlike those proposed in Leo:2018odn, match consistently the CDM prediction.
@@
We find that, under the assumption that the star formation rate (SFR) is proportional to the halo growth rate, the observations of the UV luminosity function up to high $z$ indicate two interesting effects that change the luminosity function. First, at $z>10$, either the stars need to be more luminous and/or the SFR needs to be higher. This transition is relatively sharp and might be caused by a top-heavy Pop III-dominated star population. We find that the evidence for this effect above the $95\%$ CL. In addition, the SFR needs to be further enhanced at $z>15$. This can be achieved by reducing the feedback from supernovae (SNe) and active galactic nuclei (AGNs) on the SFR. We find that to explain the $z>15$ observations, these feedbacks need to be significantly reduced to match the UV luminosity function, and at $z\simeq 25$ the production of stars should be almost free from feedback effects. In agreement with Yung:2025ttv, we find that, with these changes in the stellar population and feedback effects, the JWST data can be explained within the CDM model.
@@
An important focus of this work is whether different DM scenarios require different star formation physics to fit high-redshift data. Interestingly, we find very little sensitivity to DM properties that differ from CDM, with only a small change in star formation rates favoured for WDM, and no change for FDM.
@@
When marginalizing over the SFR parametrization, we find a lower bound of $m_{\rm FDM}>5.6\times 10^{-22}{\rm eV}$ for the case of FDM and $m_{\rm WDM}>1.5\,{\rm keV}$ for the case of WDM. The new JWST data, although pushing the observations to much earlier galaxies, are still not precise enough to put much stronger bounds than the HST data, as also found in Winch:2024mrt. We also find that enhanced small-scale matter perturbations are not preferred over changes in the star formation due to the presence of Pop-III stars and changes in the feedback in the pre-reionisation era. We are able to put bounds on the scale of white-noise enhanced perturbations at $k_{c}>25\,{\rm Mpc}^{-1}$, excluding, for example, the possibility of axion miniclusters produced for $m_{a}<6.6\times 10^{-17}\,{\rm eV}$ or heavy primordial black holes (PBHs) that constitute the fraction $f_{\rm PBH}>\max[105M_{\odot}/m_{\rm PBH},10^{-4}(m_{\rm PBH}/10^{4}M_{\odot})^{-0.09}]$ of DM. In a forthcoming paper, we will leverage these findings to compute the growth of SMBHs in non-CDM scenarios and obtain, for the first time, even more competitive bounds on deviations from CDM using the measurements of high-$z$ SMBHs by JWST.
@@
Throughout this work, we use the Planck 2018 CMB best-fit values for the cosmological parameters (Planck:2018vyg): $\Omega_{\rm M}=0.315$, $\Omega_{\rm B}=0.0493$, $z_{\rm eq}=3402.0$, $\sigma_{8}=0.811$, $h=0.674$, $T_{0}=2.7255\,{\rm K}$ and $n_{s}=0.965$.
@@ 2 Halo mass function and growth rate
Figure 1: Left panel: The unconditional first-crossing distribution for walks starting from $S=0$, $\delta=0$. The histogram shows the results of random walks, the solid black and dashed orange curves show the ansatz (3) with $q=0.80$ and $q=0.707$, respectively, and the dotted grey curve shows the spherical collapse result. Other panels: The conditional first-crossing distribution at $z>z_{0}$ for walks starting from $S_{0}>0$ and $\delta_{\rm ell}(z_{0},S_{0})>0$ as indicated in the plot. The histogram again shows the results of random walks, the solid black curves show the ansatz (LABEL:eq:ST2), the dashed orange curves show the ansatz (3) with $\delta_{\rm sp}(z)\to\delta_{\rm sp}(z)-\delta_{\rm sp}(z_{0})$ and $S\to S-S_{0}$, and the dotted gray curve shows the spherical collapse result.
In the excursion-set formalism (Bond:1990iw), the conditional and unconditional HMFs are obtained from the first-threshold-crossing distributions of uncorrelated random walks in the $(S,\delta)$ plane, where $S=\sigma(M)^{2}$ denotes the variance of matter perturbations at mass scale $M$ and $\delta$ is the matter density contrast around a given spatial point. The unconditional mass function is obtained from the first-crossing distribution $p_{\rm FC}(S,z)$ of walks starting from $(0,0)$:
@@
For spherical collapse in a universe with CDM, the density contrast threshold is $\delta_{\rm sp}(z)=3(3\pi/2)^{2/3}/(5D_{\rm g}(z))$ (Dodelson:2003ft), where the CDM linear growth function $D_{\rm g}(z)$ does not depend on $S$ and the first-crossing distribution $p_{\rm FC}(S,z)$ can be computed analytically. In the extension to include ellipsoidal collapse, the threshold becomes $S$-dependent:
@@
with $a=0.707$ (Sheth:1999su), and $p_{\rm FC}(S,z)$ cannot be calculated analytically. The standard Sheth-Tormen ansatz for the first-crossing distribution with ellipsoidal collapse is (Sheth:1999mn)
@@
where the normalization factor $A$ is given by $A=[1+2^{-p}\Gamma(1/2-p)/\sqrt{\pi}]^{-1}$, $p=0.3$ and q=a=0.707q=a=0.707.111The spherical collapse model corresponds to $p=0$, $q=1$ and $A=1/2$. In the left panel of Fig. 1, we show that random walks prefer a slightly higher value of $q=0.80$. The histograms shown in Fig. 1 are obtained by generating $2\times 10^{5}$ random walks with step size $\Delta\nu/\nu=10^{-4}$. The Seth-Tormen halo mass function has been found to match well the results of numerical simulations (see, e.g., 2003MNRAS.346..565R; Zheng:2023myp).
For the conditional mass function, we study random walks starting from $S_{0}>0$ and $\delta_{0}=\delta_{\rm ell}(S_{0},z_{0})>0$. The simplest ansatz for the first-crossing distribution $p_{\rm FC}(S,z|S_{0},z_{0})$, as considered, e.g., by Sheth:2001dp, is obtained by the shift $\delta_{\rm sp}(z)\to\delta_{\rm sp}(z)-\delta_{\rm sp}(z_{0})$ and $S\to S-S_{0}$ in (3). As seen in Fig. 1, comparing the orange dashed curves with the histograms, we find that this ansatz does not reproduce well the results of the random walks. A better estimate, shown by the solid black curves in Fig. 1, is given by
@@
In the limit $z\to z_{0}$, our formula (LABEL:eq:ST2) agrees well with the analytical estimate derived by Zhang:2008dx. However, our expression also matches the numerical results for large redshift differences $z-z_{0}\gg 0$.  [^2] We note that the ellipsoidal collapse thresholds at different redshifts can intersect. This occurs, however, only at very small halo masses, $M<10^{4}M_{\odot}$, which are not relevant for our study.
@@
Unlike the estimate derived by Correa:2014xma, our estimate (7) accounts for the correction from ellipsoidal collapse. Moreover, our estimate is based on the expected increase in the halo mass, while the estimate of Correa:2014xma is based on finding the expected mass of the main progenitor. However, we find that these estimates agree within a factor two across a broad range of halo masses and redshifts.
@@ 3.1 Suppressed small-scale structures
where $\mathcal{P}_{\rm CDM}(k)$ denotes the CDM matter power spectrum that we compute with the transfer function from Eisenstein:1997ik.
In the case of FDM, the speed of sound is modified by quantum pressure, so that $c_{s}\sim k/(4a^{2}m_{\rm FDM}^{2})$. Consequently, the Jeans scale is non-zero (see e.g. Marsh:2015xka),
@@
and perturbations at small scales $k\gtrsim k_{\rm J}$ are suppressed. This suppression can be estimated using an effective fluid approximation that is accurately matched to the solution of the Klein-Gordon equation (Hu:2000ke; Passaglia:2022bcr). We use the fitting formula provided by Passaglia:2022bcr:
@@
In the case of WDM, a thermal production mechanism is typically assumed, and the particles are lighter than typical CDM particles, remaining relativistic for a longer time. This implies that the free streaming length of WDM particles can be long, $\lambda_{\rm fs}\sim(m_{\rm WDM}/{\rm keV})^{-4/3}{\rm Mpc}$. In order to compute the resultant damping of the small-scale structures, it is necessary to track accurately the transition between relativistic and non-relativistic behaviour by solving the Boltzmann equations. This yields the fit (Bode:2000gq; Hansen:2001zv; Viel:2005qj)